OpenStudy (anonymous):

divide (c^6+4c^4+2c^2) by c^2

OpenStudy (anonymous):

\(\large \frac{c^6+4c^4+2c^2}{c^2}=\frac{c^6}{c^2}+ \frac{4c^4}{c^2}+ \frac{2c^2}{c^2} \) = ??? simplify..